# Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex Optimization

### v2

@@ 4.2 Decentralized generalized SGD with momentum
Furthermore, from the definition of $f$ in Assumption 4.1(1), we attain that $\Phi({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\mathcal{D}_{f}({\bm{x}})\end{smallmatrix}\right]$. Then together with [36, Proposition 4.5] and the path-differentiability of $f$, we can conclude that $\psi({\bm{x}},{\bm{y}})=f({\bm{x}})+\frac{1}{\tau}\phi({\bm{y}})$ is the Lyapunov function of the differential inclusion $\left(\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t},\frac{\mathrm{d}{\bm{y}}}{\mathrm{d}t}\right)\in-\Phi({\bm{x}},{\bm{y}})$, and admits $\{({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ as its stable set. Together with the definability of $f$ and $\mathcal{D}_{f}$, we can conclude that the set $\{\psi({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ is finite in $\mathbb{R}$. Therefore, Assumption 3.3 holds for Algorithm 2.
Finally, from Theorem 3.9, we can conclude that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-stationary point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof. ∎
@@ 4.3 Decentralized SGD with gradient-tracking technique
The decentralized SGD method with gradient-tracking technique (DSGD-T) is another widely employed stochastic (sub)gradient-based method in decentralized learning. Algorithm 3 exhibits the detailed implementations of DSGD-T in minimizing nonsmooth path-differentiable functions. Different from the DSGD in Algorithm 1, DSGD-T introduces a sequence of auxiliary variables to track the average of local subgradients over all the agents. To the best of our knowledge, all the existing works focus on analyzing the convergence of DSGD-T with Clarke regular objective functions, and the convergence of DSGD-T for nonsmooth nonconvex functions remains unexplored.
In this subsection, we prove that DSGD-T can be enclosed by our proposed framework (DSM). Therefore, based on our theoretical results exhibited in Theorem 3.9, we can establish the global convergence of DSGD-T in the minimization of nonsmooth path-differentiable functions.
@@
1:Initial point ${\bm{x}}_{0}=[{\bm{x}}_{1,0},\ldots,{\bm{x}}_{d,0}]$, initial descent direction ${\bm{v}}_{0}=[{\bm{v}}_{1,0},\ldots,{\bm{v}}_{d,0}]$ satisfying ${\bm{v}}_{0}{\bm{1}}_{d}=\sum_{i=1}^{d}{\bm{d}}_{i,0}$, where ${\bm{d}}_{i,0}\in\frac{1}{|\mathcal{B}_{i,0}|}\sum_{l\in\mathcal{B}_{i,0}}D_{F(\cdot;{\bm{s}}_{l})}({\bm{x}}_{i,0}),$ $\mathcal{B}_{i,0}$ is a mini-batch of $\mathcal{S}_{i}$, and a mixing matrix ${\bm{W}}$.
@@
8:         Compute ${\bm{d}}_{i,k+1}\in\frac{1}{|\mathcal{B}_{i,k+1}|}\sum_{l\in\mathcal{B}_{i,k+1}}D_{F(\cdot;{\bm{s}}_{l})}({\bm{x}})$;
@@
In the rest of this subsection, we denote $,{\bm{V}}_{k}=[{\bm{v}}_{1,k},$$\ldots,{\bm{v}}_{d,k}]$ at $k$-th iteration in Algorithm 3. Assumption 4.3 and 4.1 are also adopted for DSGD-T. In the following, we condense the assertion that DSGD-T can fit in our framework (DSM), along with convergence analysis, into a single theorem.
@@
Suppose Assumption 4.1 and 4.3 hold for Algorithm 3. Then the iterative formula of $\{{\bm{X}}_{k}\}$ is a specific form of our proposed general framework (DSM). Moreover, for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-stationary point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
@@
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. By setting ${\bm{Z}}_{k}={\bm{X}}_{k}$, ${\bm{H}}_{k}=\mathbb{E}[{\bm{V}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}={\bm{V}}_{k}-{\bm{H}}_{k}$, we can conclude that (4.5) coincides with the framework (DSM).
Then we aim to check the validity of Assumption 3.1. From the uniform boundedness of $\{{\bm{X}}_{k}\}$ in Assumption 4.3(1), we can conclude that there exists $M_{D}>0$ such that $\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\leq M_{D}$ almost surely. Then by the update scheme of $\{{\bm{V}}_{k}\}$ in (4.5), it holds that
${\bm{V}}_{k+1}={}$ ${\bm{D}}_{k+1}{\bm{W}}+\sum_{i=1}^{k}{\bm{D}}_{i}({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k+1-i}$
$-{\bm{D}}_{0}{\bm{W}}^{k+1}+{\bm{V}}_{0}{\bm{W}}^{k+1}.$
Then from Proposition 2.2, there exists a constant $\alpha\in(0,1)$, such that $\left\|({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k}\right\|\leq 2\alpha^{k}$ holds for any $k\geq 0$. Therefore, almost surely, it holds that
$\sup_{k\geq 0}\left\|{\bm{V}}_{k}\right\|$
[…]
$-{\bm{D}}_{0}{\bm{W}}^{k+1}+{\bm{V}}_{0}{\bm{W}}^{k+1}\|$
@@
Therefore, we can conclude that $\{{\bm{V}}_{k}\}$ is uniformly bounded almost surely, hence verifying the validity of Assumption 3.1(1).
Moreover, the uniform boundedness of $\{\Xi_{k}\}$ and [30, Remark 1.5] illustrates that Assumption 3.1(2) holds almost surely for Algorithm 3. In addition, followed by $\mathbb{E}[{\bm{d}}_{i,k}|\mathcal{F}_{k}]\in\mathbb{E}_{\mathcal{B}_{i,k}}[D_{\mathcal{B}_{i,k}}({\bm{x}}_{i,k})]=D_{f_{i}}({\bm{x}}_{i,k})$, it is easy to verify that
$\frac{1}{d}{\bm{H}}_{k}{\bm{1}}_{d}$ $=\mathbb{E}[\frac{1}{d}{\bm{V}}_{k}{\bm{1}}_{d}|\mathcal{F}_{k}]=\mathbb{E}[\frac{1}{d}{\bm{D}}_{k}{\bm{1}}_{d}|\mathcal{F}_{k}]$
$\in\frac{1}{d}\sum_{i=1}^{d}\mathcal{D}_{f_{i}}({\bm{x}}_{i,k}),\forall k\in\mathbb{N}.$
Therefore, by choosing $\Phi_{i}=\mathcal{D}_{f_{i}}$ for all $i\in[d]$, we verify the validity of Assumption 3.1(3).
Furthermore, according to [28, Section 6] and Remark 4.2, we can assert that $f$ is the Lyapunov function of the differential inclusion $\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-\mathcal{D}_{f}({\bm{x}})$, and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. In addition, the definability of $f$ and $\mathcal{D}_{f}$ illustrates that $\{f({\bm{x}}):{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we have verified the validity of Assumption 3.3 with $\Phi=\mathcal{D}_{f}$, $\psi=f$, and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$.
Finally, from Theorem 3.9, we can conclude that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-stationary point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof.
@@ 5 Numerical Experiments
Figure 1: Numerical performance comparison of DSGD, DSGD-M and DSignSGD in training ResNet50 on CIFAR-100 dataset.
@@ 5.1 Experiment settings
We conduct a series of image classification tasks with CIFAR-10 and CIFAR-100 datasets [54], which both consist of $50000$ training samples and $10000$ test samples. ResNet-50 [55] is employed as our nonsmooth neural network. The decentralized network topology is set as ring-structure with $8$ agents, and the mixing matrix is chosen as the Metropolis weight matrix [46]. In each experiment, we divide the training dataset into 8 equal parts, with each part serving as the local dataset for each agent. During any epoch of training, the local dataset of each agent itself remains unchanged, but the mini-batch is reshuffled to ensure compliance with the format of our problem (DOP) and the randomness requirements of all algorithms in Section 4.
Moreover, we set the number of epochs as $200$ and choose the batch size as $128$. The strategy for selecting step-size is to decrease it at the $60$-th, $120$-th and $160$-th epochs, with a decaying factor of $0.2$.

### v3

@@ 4.2 Decentralized generalized SGD with momentum
Furthermore, from the definition of $\Phi_{i}$ and (1.5), we attain that $\Phi({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau D_{f}({\bm{x}})\end{smallmatrix}\right]$. Then together with [43, Proposition 4.5] and the path-differentiability of $f$, we can conclude that $\psi({\bm{x}},{\bm{y}})=f({\bm{x}})+\frac{1}{\tau}\phi({\bm{y}})$ is the Lyapunov function of the differential inclusion $\left(\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t},\frac{\mathrm{d}{\bm{y}}}{\mathrm{d}t}\right)\in-\Phi({\bm{x}},{\bm{y}})$, and admits $\{({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ as its stable set. Leveraging the definability of $f$ and $D_{f}$, it follows that the set $\{\psi({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ is finite in $\mathbb{R}$. Therefore, Assumption 3.1 holds for Algorithm 2.
We set $\mathcal{U}_{i,j}({\bm{x}},{\bm{y}}):=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\frac{1}{|\mathcal{B}_{i,j}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,j}}D_{F_{i,l}}({\bm{x}})\end{smallmatrix}\right]$, where $\mathcal{B}_{i,j}$ for $10\leq j\leq N-$ 1 are subsets of equal size, and set $\rho_{k}:=\eta_{k}\|{\bm{T}}_{k}\|+2\sum_{j=0}^{k-1}\lambda_{2}^{k-1-j}\eta_{j}\|{\bm{T}}_{j}\|$. Then (4.4) is a special form of (DSM) with (3.8). Moreover, Assumption 4.2 implies that Assumption 3.8 is satisfied for case (1). With the choice of $\xi_{i,k}:=\{\textit{Draw out }\mathcal{B}_{i,k}\textit{ from agent }i\}$, and
$\chi_{i}({\bm{x}},{\bm{y}},\xi_{i,k})\in\left[\begin{smallmatrix}0\\ \frac{-\tau}{|\mathcal{B}_{i,k}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k}}D_{F_{i,l}}({\bm{x}},{\bm{y}})-D_{f_{i}}({\bm{x}},{\bm{y}})\end{smallmatrix}\right],$
(4.4) is a particular form of (DSM) with (3.18), and Assumption 3.12 is satisfied for case (2). Plugging Lemma 3.6 and Theorem 3.14, we can prove that $\delta_{k}\to 0$, as $k\to\infty$ with arbitrary high probability  [^2] Notice that Theorem 3.14 does not require $\delta_{k}$ to diminish.. Employing Theorem 3.11 and 3.15, we can derive the results in Theorem 4.7.
Theorem 4.8 illustrates how Algorithm 2 avoids spurious critical points, whose proof is presented in Appendix A for simplicity.
Theorem 4.8.
Suppose Assumption 4.1 holds and ${\bm{W}}$ is a non-singular mixing matrix. $\mathcal{X}_{0}$ is a compact set of $\mathbb{R}^{n}$, sequence $\{{\bm{X}}_{k}\}$ is generated by Algorithm 2, and auxiliary function $\phi$ is convex which admits a unique minimizer at ${\bm{0}}.$
  1. (1)
    When Assumption 4.2 holds, there exists $\alpha_{c}>0$, a full-measure subset $\mathcal{S}$ of $(0,\alpha_{c})$, and a full-measure subset $\mathcal{K}$ of $\mathcal{X}_{0}\times\mathbb{R}^{n}$, such that for any $c\in\mathcal{S}$ and $({\bm{x}}_{0},{\bm{y}}_{0})\in\mathcal{K}$, it holds that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a Clarke-critical point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges for any $i\in[d]$.
  2. (2)
    When Assumption 4.3 holds, for any $\varepsilon>0$, there exists $\alpha_{c}>0$, a full-measure subset $\mathcal{S}$ of $(0,\alpha_{c})$, and a full-measure subset $\mathcal{K}$ of $\mathcal{X}_{0}\times\mathbb{R}^{n}$, such that for any $c\in\mathcal{S}$ and $({\bm{x}}_{0},{\bm{y}}_{0})\in\mathcal{K}$,
    $\mathbb{P}\left(\lim_{k\to\infty}\mathrm{dist}({\bm{X}}_{k},\{{\bm{X}}\in\mathbb{R}^{n\times d}:{\bm{X}}={\bm{x}}{\bm{1}}^{\top},{\bm{0}}\in\partial f({\bm{x}})\})=0\right)\geq 1-\varepsilon$
    and
    $\mathbb{P}(f({\bm{x}}_{i,k})\textit{ converges })\geq 1-\varepsilon,\textit{ }\forall i\in[d].$
@@ 4.3 Decentralized SGD with gradient-tracking technique
The decentralized SGD method with gradient-tracking technique (DSGD-T) is another widely employed stochastic (sub)gradient-based method in decentralized learning. Algorithm 3 exhibits the detailed implementations of DSGD-T in minimizing nonsmooth path-differentiable functions. Different from the DSGD in Algorithm 1, DSGD-T introduces a sequence of auxiliary variables to track the average of local subgradients over all the agents. To the best of our knowledge, all the existing works focus on analyzing the convergence of DSGD-T with Clarke regular objective functions, and the convergence of DSGD-T for nonsmooth functions without Clarke regularity remains unexplored.
In this subsection, we prove that DSGD-T can be enclosed by our proposed framework (DSM) with (3.8) or (3.18). Therefore, based on our theoretical results exhibited in Theorem 3.11 and 3.15, we can establish the global convergence of DSGD-T in the minimization of nonsmooth path-differentiable functions.
@@
1:Initial point ${\bm{x}}_{0}\in\mathbb{R}^{n}$, initial descent direction ${\bm{V}}_{0}=[{\bm{v}}_{1,0},\ldots,{\bm{v}}_{d,0}]$ satisfying ${\bm{V}}_{0}{\bm{1}}_{d}=\sum_{i=1}^{d}{\bm{d}}_{i,0}$, where ${\bm{d}}_{i,0}\in\frac{1}{|\mathcal{B}_{i,0}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,0}}D_{F_{i,l}}({\bm{x}}_{i,0}),$ $\mathcal{B}_{i,0}$ is a mini-batch of $\mathcal{S}_{i}$, and a mixing matrix ${\bm{W}}$.
@@
8:         Compute ${\bm{d}}_{i,k+1}\in\frac{1}{|\mathcal{B}_{i,k+1}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k+1}}D_{F_{i,l}}({\bm{x}}_{i,k+1})$;
@@
In the rest of this subsection, we denote $,{\bm{V}}_{k}=[{\bm{v}}_{1,k},$$\ldots,{\bm{v}}_{d,k}]$ at $k$-th iteration in Algorithm 3. We condense the assertion that DSGD-T can fit in our framework (DSM) with different noise settings, along with convergence analysis, into a single theorem.
@@
Suppose Assumption 4.1 holds. Let $\{{\bm{X}}_{k}\}$ be the sequence generated by Algorithm 3.
  1. (1)
    When Assumption 4.2 holds, there exists $\alpha_{c}>0$, for any $c\in(0,\alpha_{c})$, it holds that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $D_{f}$-critical point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges for any $i\in[d]$.
  2. (2)
    When Assumption 4.3 holds, for any $\varepsilon>0$, there exists $\alpha_{c}>0$, for any $c\in(0,\alpha_{c})$,
    $\mathbb{P}\left(\lim_{k\to\infty}\mathrm{dist}({\bm{X}}_{k},\{{\bm{X}}\in\mathbb{R}^{n\times d}:{\bm{X}}={\bm{x}}{\bm{1}}^{\top},{\bm{0}}\in D_{f}({\bm{x}})\})=0\right)\geq 1-\varepsilon$
    and
    $\mathbb{P}(f({\bm{x}}_{i,k})\textit{ converges })\geq 1-\varepsilon,\textit{ }\forall i\in[d].$
@@
Let the filtration $\mathcal{F}_{k}:=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. By setting ${\bm{Z}}_{k}:={\bm{X}}_{k}$, ${\bm{H}}_{k}:=\mathbb{E}[{\bm{V}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}:={\bm{V}}_{k}-{\bm{H}}_{k}$, we can conclude that (4.5) coincides with the framework (DSM).
Then we aim to check the validity of Assumption 3.1. Due to the local boundedness of any conservative field, there exists $M_{D}>0$ such that $\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\leq M_{D}$ almost surely when $\{{\bm{X}}_{k}\}$ is restricted in some bounded set. Then by the update scheme of $\{{\bm{V}}_{k}\}$ in (4.5), it holds that
${\bm{V}}_{k+1}={\bm{D}}_{k+1}{\bm{W}}+\sum_{i=1}^{k}{\bm{D}}_{i}({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k+1-i}-{\bm{D}}_{0}{\bm{W}}^{k+1}+{\bm{V}}_{0}{\bm{W}}^{k+1}.$
Then from Corollary 2.2, there exists a constant $\alpha\in(0,1)$, such that $\left\|({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k}\right\|\leq 2\alpha^{k}$ holds for any $k\geq 0$. Therefore, almost surely, it holds that
$\sup_{k\geq 0}\left\|{\bm{V}}_{k}\right\|\leq{}$ $\sup_{k\geq 0}\|{\bm{D}}_{k+1}{\bm{W}}+\sum_{i=1}^{k}{\bm{D}}_{i}({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k+1-i}-{\bm{D}}_{0}{\bm{W}}^{k+1}+{\bm{V}}_{0}{\bm{W}}^{k+1}\|$
@@
Therefore, we can conclude that $\{{\bm{V}}_{k}\}$ is uniformly bounded almost surely when $\{{\bm{X}}_{k}\}$ is bounded, hence verifying the validity of Assumption 3.1(2).
Assumption 3.1(1) follows quickly from Assumption 4.2 or 4.3. (1.5) holds vacuously with sufficiently large $\delta_{k}$ and $\Phi_{i}:=D_{f_{i}}$ for case (1). Moreover for case (2), followed by
$\mathbb{E}[{\bm{d}}_{i,k}|\mathcal{F}_{k}]\in\mathbb{E}_{\mathcal{B}_{i,k}}[\frac{1}{|\mathcal{B}_{i,k}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k}}D_{F_{i,l}}({\bm{x}}_{i,k})]=D_{f_{i}}({\bm{x}}_{i,k}),$ (4.6)
it is easy to deduce that
$\frac{1}{d}{\bm{H}}_{k}{\bm{1}}_{d}=\mathbb{E}[\frac{1}{d}{\bm{V}}_{k}{\bm{1}}_{d}|\mathcal{F}_{k}]=\mathbb{E}[\frac{1}{d}{\bm{D}}_{k}{\bm{1}}_{d}|\mathcal{F}_{k}]\in\frac{1}{d}\sum_{i=1}^{d}D_{f_{i}}({\bm{x}}_{i,k}),\forall k\in\mathbb{N}.$
Therefore, by choosing $\delta_{k}:=0$ and $\Phi_{i}:=D_{f_{i}}$ for all $i\in[d]$, we verify the validity of (1.5).
Furthermore, according to [36, Section 6], we can assert that $f$ is the Lyapunov function of the differential inclusion $\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-D_{f}({\bm{x}})$, and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$. In addition, the definability of $f$ and $D_{f}$ illustrates that $\{f({\bm{x}}):{\bm{0}}\in D_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we have verified the validity of Assumption 3.1 with $\Phi:=D_{f}$, $\psi:=f$, and $\mathcal{A}:=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$.
Following the same construction of $\mathcal{U}_{i,k}({\bm{x}})$, $\rho_{k}$, $\xi_{i,k}$ and $\chi_{i}$ as in the proof of Theorem 4.4, we could verify Assumption 3.8 and 3.12 hold for case (1) and case (2), respectively. Henceforth, the update scheme (4.5) is a variant of (DSM) with either (3.8) or (3.18). Utilizing our theoretical results in Theorem 3.11 and 3.15, we completes the proof.
@@ 5 Numerical Experiments
Figure 1: Numerical performance comparison of DSGD, DSGD-M, and DSignSGD in training ResNet50 on CIFAR-100 dataset.
@@ 5.1 Experiment settings
We conduct a series of image classification tasks with CIFAR-10 and CIFAR-100 datasets [67], which both consist of $50000$ training samples and $10000$ test samples. ResNet-50 [68] is employed as our nonsmooth neural network. The decentralized network topology is set as ring-structure with $8$ agents, and the mixing matrix is chosen as the Metropolis weight matrix [54]. In each experiment, we divide the training dataset into 8 equal parts, with each part serving as the local dataset for each agent. During any epoch of training, the local dataset of each agent itself remains unchanged, while the mini-batch is selected by random reshuffling to ensure compliance with the sampling scheme in Assumption 4.2.
Moreover, we set the number of epochs as $200$ and choose the batch size as $128$. The strategy for selecting step-size is designed to reduce it three times, specifically at the 60-th, 120-th, and 160-th epochs, with a decay factor of 0.2.